# PathoDuet: Foundation Models for Pathological Slide Analysis of H&E and IHC Stains

### v1

@@ Abstract
Large amounts of digitized histopathological data display a promising future for developing pathological foundation models via self-supervised learning methods. Foundation models pretrained with these methods serve as a good basis for downstream tasks. However, the gap between natural and histopathological images hinders the direct application of existing methods. In this work, we present PathoDuet, a series of pretrained models on histopathological images, and a new self-supervised learning framework in histopathology. The framework is featured by a newly-introduced pretext token and later task raisers to explicitly utilize certain relations between images, like multiple magnifications and multiple stains. Based on this, two pretext tasks, cross-scale positioning and cross-stain transferring, are designed to pretrain the model on Hematoxylin and Eosin (H&E) images and transfer the model to immunohistochemistry (IHC) images, respectively. To validate the efficacy of our models, we evaluate the performance over a wide variety of downstream tasks, including patch-level colorectal cancer subtyping and whole slide image (WSI)-level classification in H&E field, together with expression level prediction of IHC marker and tumor identification in IHC field. The experimental results show the superiority of our models over most tasks and the efficacy of proposed pretext tasks. The codes and models are available at https://github.com/openmedlab/PathoDuet.
Keywords Foundation Model  $\cdot$ Pathological Image  $\cdot$ H&E and IHC
@@ 1 Introduction
The histologic assessment stands as the gold standard for diagnosing specific cancers, predominantly depending on the expertise of pathologists. The assessment is mainly based on the analysis of Hematoxylin and Eosin (H&E) stained slides, offering fundamental structural information. Pathologists may further augment their conclusions by utilizing functional stains such as Immunohistochemistry (IHC) to provide additional diagnostic insights. As technology continues to advance, digital scanners with high throughput have revolutionized the acquisition of pathological data. Despite large amounts of data, the integration of deep learning techniques into diagnostic processes has been progressing at a relatively measured pace. This can be attributed, in part, to the limited amount of labeled data for certain tasks. Unlike the annotation of natural images, the annotation process for pathological images demands expertise, rendering it resource-intensive and time-consuming. To address this challenge, foundation models emerge as a prospective solution. These models typically exploit the potential of unlabeled data, facilitating efficient transferring to downstream tasks with reduced dependency on labeled data.
@@
Existing foundation models mainly rely on self-supervised learning (SSL) methodologies. The essence of SSL involves the generation of supervised signals directly from the data itself. This process is often called the pretext task (Jing and Tian, 2020). As a dominant branch of SSL methods, contrastive learning (CL) has attracted significant attention (He et al., 2020; Chen et al., 2020; Oquab et al., 2023). In general, CL focuses on exploiting image similarity as a means to discern and categorize images concerning others. Another branch, featured by masked autoencoders (He et al., 2022), utilizes image generation to boost models’ understanding. Compared with generative SSL methods, CL has better performance when transferred to discriminative tasks (Shekhar et al., 2023). As a result, CL is preferred in this work given the fact that quite a few pathological tasks are highly related to identification. However, the direct application of CL methods designed for natural images to histopathological images requires careful consideration. CL posits that the majority of images should possess semantic uniqueness. A routine is to contract different views of the same image and to separate different images in semantic space. Pathological whole slide images (WSI) are yet cropped into smaller patches to fit in the input size requirements of most models, restricting cropped patches to exhibit semantic distinctiveness from neighboring patches. Tough isolation of these patches may cause over-fragmentation of semantic space, thus affecting the performance of models. The conflict, consequently, requires a special design of contrastive pretext tasks concerning the essence of histopathological images. In seeking insights for the strategy of tailoring, inspiration can be drawn from the analogous process of pathological evaluation.
@@
A typical characteristic of pathologists’ working methodology is their habitual practice of zooming in and out during the examination process. Initially, they employ a low magnification level to screen overall structures and tissues, identifying regions of interest that require closer inspection. Subsequently, at a higher magnification level, pathologists analyze individual cells or clusters of cells, refining their understanding and classification of the identified regions. To simulate the zooming in and out operations performed by pathologists, we define one of our pretext tasks as a “cross-scale positioning" task, leveraging the large-scale public H&E datasets to develop a foundation model for H&E stain. In which, besides the commonly used two branches of different augmented image views in CL, we add a branch that learns the representation of a patch from its neighboring regions. This manipulation enables patches to be understood from a broader perspective, thereby alleviating the conflict between CL’s requirement on semantic division and pathological patches’ concentration.
@@
Additionally, pathologists often utilize additional functional slides for a more comprehensive diagnosis. Notably, IHC markers are frequently employed, offering valuable insights into subtyping cancers. However, the effective interpretation of IHC slides cannot be derived without H&E stained slides. The H&E slides serve as a fundamental reference, providing essential contextual information and structural details to complement the specific molecular information gleaned from IHC slides. Therefore, an ideal foundation model for IHC stain should be able to assess IHC images according to markers’ expression levels and align with H&E models in the semantic space in the aspect of the tissue structure. With the limited publicly available IHC data, we exploit the trained H&E foundation model and introduce the “cross-stain transferring" pretext task, to deepen comprehension of pathological images stained in a different way. Specifically, we align the IHC representation with the IHC-style transferred H&E representation via a transferer drawing on the adaptive instance normalization, AdaIN (Huang and Belongie, 2017). This alignment injects structural information that is readily accessible in H&E images, as well as preserving diagnostic information rooted in IHC images.
@@
Figure 1: An overview of PathoDuet. Left: two pretext tasks, cross-scale positioning and cross-stain transferring, are designed to develop H&E and IHC models. Right: a series of downstream tasks, covering both H&E and IHC ones, are used to evaluate models’ performance in application.
@@
A pretext token mechanism is introduced to unify the two proposed tasks. Both tasks require an auxiliary input in a different form, i.e., a much smaller patch or a staining hint. Contrary to designing a separate network to handle the additional input, an extra token with the auxiliary information is fed into the Vision Transformer (ViT) model (Dosovitskiy et al., 2020) throughout the network training, and thereby combining the cross-scale or cross-stain information with original representation. The relation is subsequently exploited by a delicately designed module, termed a task raiser, to explicitly associate the two forms. This mechanism enriches the model’s capacity to discover and leverage intrinsic correlation between tasks and staining modalities in a lightweight way.
@@
The whole framework and ensuing models are collectively denoted as PathoDuet. This reflects the dual functionality of the framework, i.e., it offers two distinct strategies for developing foundation models: 1) by exploiting a shifted view of scales to discover broader semantic space, e.g., comprehending patches further from surrounding regions, 2) by progressing on the basis of other closely related and already exploited modalities, e.g., learning representation of IHC images from H&E model. The efficacy of our proposed methods is validated across a wide range of downstream tasks, covering both H&E and IHC images. These tasks range from classifying cancer tissue types at the patch level, categorizing WSIs, and identifying cancer cells within IHC images, to assessing the expression levels of IHC markers. Our contributions are summarized as:
@@ 2.1 Self-supervised learning
Pretext tasks can be various. Some tasks aim at predicting properties of images, like rotation angles (Gidaris et al., 2018), flipping (Srinidhi et al., 2022), etc. Some design small problems that help to learn features of images, like jigsaw puzzle solving (Noroozi and Favaro, 2016) and relative position prediction of sub-regions (Doersch et al., 2015). Some exploit the process of image generation, including image colorization (Zhang et al., 2016), super-resolution (Ledig et al., 2017), inpainting (Pathak et al., 2016), and exploiting generative adversarial networks (GANs) (Zhu et al., 2017). Besides, He et al. propose masked autoencoder (MAE) (He et al., 2022) that leads recent research into masked image modeling.
Among all these pretext tasks, similarity-based CL task has demonstrated their superiority, because they focus on invariant features instead of covariant ones. CL-based methods pull together the latent representations of similar images while keeping those of dissimilar images away from each other. These methods often regard different augmented views from the same image as a similar (positive) pair, and those from different images as dissimilar (negative) pairs. Therefore, CL-based methods typically contain two network branches. MoCo v1 (He et al., 2020) is a basic one with a symmetric structure, a momentum-update mechanism, a memory bank, and an InfoNCE loss that considers positive samples and negative samples together. SimCLR (Chen et al., 2020) adds a simple multi-layer perceptron (MLP) called projector after each backbone encoder and updates two branches concurrently without a memory bank. BYOL (Grill et al., 2020) further adds another MLP, predictor, after one of the branches (the online branch), while updating the other target branch with momentum, and changing the loss to cosine similarity loss which does not involve negative samples. SimSiam (Chen and He, 2021) proposes a similar structure of BYOL but in a symmetric way with its gradient stop technique to avoid collapse. In MoCo v3 (Chen et al., 2021), it adopts the BYOL’s structure but continues using the InfoNCE loss. DINO (Caron et al., 2021) and its later version DINO v2 (Oquab et al., 2023) integrate previous works and achieve another state-of-the-art performance, thereby attracting applications to other fields.
@@ 2.2 Self-supervised learning in digital pathology
With the development of SSL frameworks in computer vision, the concept and some existing methods have been migrated to histopathology. Besides traditional pretext tasks, some works design specific pretext tasks in digital pathology. Magnification prediction (Sahasrabudhe et al., 2020; Koohbanani et al., 2021) and stain prediction (Koohbanani et al., 2021) are two simple tasks that exploit the characteristics of pathological images. Resolution sequence prediction (Srinidhi et al., 2022) is a dedicated task with easy supervision and remarkable performance. Others leverage the stain. RestainNet (Zhao et al., 2022), as its name suggests, simulates the process of de-staining and restaining via separated Hematoxylin and Eosin channels. Ling et al. (2023) use self-supervised methods to realize arbitrary stain transfer of pathological images from different domains. These methods though take into consideration the features of histopathological images, their performance depends on the association between the pretext tasks and the downstream tasks.
Some works are based on CL frameworks. Huang et al. (2021); Li et al. (2021); Ciga et al. (2022); Kawai et al. (2023) directly migrate traditional CL methods to medical data, while some studies take into account the uniqueness of histopathological images and propose a modified SSL framework. CTransPath (Wang et al., 2022) keeps the body of MoCo v3 but adds a pseudo positive selection mechanism to avoid false penalty of semantic similar patches and uses a hybrid CNN-Transformer as its backbone encoder. Meanwhile, Wang et al. also propose a clustering-guided contrastive learning method and the produced model RetCCL (Wang et al., 2023a). CS-CO (Yang et al., 2022) separates histopathological images into H-channel and E-channel and uses a cross-stain prediction task in the first phase and a CL-based method in the second phase. Abbet et al. (2022) take advantage of domain information. Meanwhile, a branch of works has focused on the gigapixel WSIs (Vu et al., 2023; Wang et al., 2023b; Lazard et al., 2023; Aryal and Yahyasoltani, 2023; Schirris et al., 2022). Recently, pathological models pretrained with ultra-large amounts of data have shown their superiority. Chen et al. (2023) collect 100,000 slides to pretrain UNI with DINO v2 framework, and Virchow (Vorontsov et al., 2023) with 1.5 million slides. Besides, some works look into multi-modal methods like visual-language learning and pretrain models on large-scale image-text datasets (Huang et al., 2023; Pisula and Bozek, 2022; Lu et al., 2023; Zhang et al., 2023), but the focus is more on leveraging strong language models.
@@ 3.1 Pretext token empowered SSL framework
The basic framework is based on a typical contrastive learning method, i.e., MoCo v3 (Chen et al., 2021). The difference starts from the input. To mimic pathologists, we need an extra input to bring the information from a different scale or stain. A typical practice to handle two input types is employing two networks respectively with a following interaction module. However, we aim at learning a single powerful encoder for the downstream tasks. Therefore, we propose the pretext token mechanism to uncover the relation within the same encoder, thus deepening understanding of pathological images like pathologists.
The mechanism starts from adding a pretext token, which is fed into the ViT together with the partitioned mini-patches’ embeddings. The token interacts with these embeddings in attention blocks and becomes an abstract association between the input pair. With the help of some simple modules, termed as task raisers, the association develops to concrete relation to perform self-supervision. The task raiser is featured by a lightweight implementation to explicitly perform the pretext task. As illustrated in the left part in Fig. 1, the task raiser is instantiated as a positioner to explore the relation between a local patch and its surrounding region in the cross-scale positioning task, or a transferer to generate IHC-style features from H&E images in the cross-stain transferring task.
In the following two subsections, we introduce in detail two pretext tasks we propose within this framework, and how pretext token and instantiated task raisers take effect in extracting information from two inputs. If not specified, we use the patch for the cropped image from a WSI and mini-patch for the output of patch embedding in ViT. Meanwhile, if the network takes only one kind of input, the final input is formulated as $[\mathbf{x};\epsilon]$, where $\mathbf{x}=[x_{1};\dots;x_{L}]\in\mathbb{R}^{L\times C}$ represents the original input after patch embedding module in ViT, and $\epsilon\in\mathbb{R}^{C}$ is a learnable placeholder of the pretext token. $L$ is the embedding length, and $C$ is the number of channels. When the network takes two inputs, the input changes to $[\mathbf{x}^{o};x^{p}]$, where $\mathbf{x}^{o}=[x^{o}_{1};\dots;x^{o}_{L}]\in\mathbb{R}^{L\times C}$ stands the embeddings of the original input, and $x^{p}\in\mathbb{R}^{C}$ those of the pretext token input, e.g., a patch in a different scale or a crop stained differently.

### v2

@@ Abstract
Large amounts of digitized histopathological data display a promising future for developing pathological foundation models via self-supervised learning methods. Foundation models pretrained with these methods serve as a good basis for downstream tasks. However, the gap between natural and histopathological images hinders the direct application of existing methods. In this work, we present PathoDuet, a series of pretrained models on histopathological images, and a new self-supervised learning framework in histopathology. The framework is featured by a newly-introduced pretext token and later task raisers to explicitly utilize certain relations between images, like multiple magnifications and multiple stains. Based on this, two pretext tasks, cross-scale positioning and cross-stain transferring, are designed to pretrain the model on Hematoxylin and Eosin (H&E) images and transfer the model to immunohistochemistry (IHC) images, respectively. To validate the efficacy of our models, we evaluate the performance over a wide variety of downstream tasks, including patch-level colorectal cancer subtyping and whole slide image (WSI)-level classification in H&E field, together with expression level prediction of IHC marker, tumor identification and slide-level qualitative analysis in IHC field. The experimental results show the superiority of our models over most tasks and the efficacy of proposed pretext tasks. The codes and models are available at https://github.com/openmedlab/PathoDuet.
keywords:
[…]
journal: Medical Image Analysis
@@ 1 Introduction
The histologic assessment stands as the gold standard for diagnosing specific cancers, predominantly depending on the expertise of pathologists. The assessment is mainly based on the analysis of Hematoxylin and Eosin (H&E) stained slides, offering fundamental structural information. Pathologists may further augment their conclusions by utilizing functional stains such as Immunohistochemistry (IHC) to provide additional diagnostic insights. As technology continues to advance, digital scanners with high throughput have revolutionized the acquisition of pathological data. Despite large amounts of data, the integration of deep learning techniques into diagnostic processes has been progressing at a relatively measured pace. This can be attributed, in part, to the limited amount of labeled data for certain tasks. Unlike the annotation of natural images, the annotation process for pathological images demands expertise, rendering it resource-intensive and time-consuming. To address this challenge, foundation models emerge as a prospective solution in many medical fields [59, 61, 32, 54, 6, 49, 48, 22, 38]. These models typically exploit the potential of unlabeled data, facilitating efficient transferring to downstream tasks with reduced dependency on labeled data.
@@
Existing foundation models mainly rely on self-supervised learning (SSL) methodologies. The essence of SSL involves the generation of supervised signals directly from the data itself. This process is often called the pretext task [23]. As a dominant branch of SSL methods, contrastive learning (CL) has attracted significant attention [18, 7, 40]. In general, CL focuses on exploiting image similarity as a means to discern and categorize images concerning others. Another branch, featured by masked autoencoders [17], utilizes image generation to boost models’ understanding. Compared with generative SSL methods, CL has better performance when transferred to discriminative tasks [46]. As a result, CL is preferred in this work given the fact that quite a few pathological tasks are highly related to identification. However, the direct application of CL methods designed for natural images to histopathological images requires careful consideration. CL posits that the majority of images should possess semantic uniqueness. A routine is to contract different views of the same image and to separate different images in semantic space. Pathological whole slide images (WSI) are yet cropped into smaller patches to fit in the input size requirements of most models, restricting cropped patches to exhibit semantic distinctiveness from neighboring patches. Tough isolation of these patches may cause over-fragmentation of semantic space, thus affecting the performance of models. The conflict, consequently, requires a special design of contrastive pretext tasks concerning the essence of histopathological images. In seeking insights for the strategy of tailoring, inspiration can be drawn from the analogous process of pathological evaluation.
@@
A typical characteristic of pathologists’ working methodology is their habitual practice of zooming in and out during the examination process. Initially, they employ a low magnification level to screen overall structures and tissues, identifying regions of interest that require closer inspection. Subsequently, at a higher magnification level, pathologists analyze individual cells or clusters of cells, refining their understanding and classification of the identified regions. To simulate the zooming in and out operations performed by pathologists, we define one of our pretext tasks as a “cross-scale positioning” task, leveraging the large-scale public H&E datasets to develop a foundation model for H&E stain. In which, besides the commonly used two branches of different augmented image views in CL, we add a branch that learns the representation of a patch from its neighboring regions. This manipulation enables patches to be understood from a broader perspective, thereby alleviating the conflict between CL’s requirement on semantic division and pathological patches’ concentration.
@@
Additionally, pathologists often utilize additional functional slides for a more comprehensive diagnosis. Notably, IHC markers are frequently employed, offering valuable insights into subtyping cancers. However, the effective interpretation of IHC slides cannot be derived without H&E stained slides. The H&E slides serve as a fundamental reference, providing essential contextual information and structural details to complement the specific molecular information gleaned from IHC slides. Therefore, an ideal foundation model for IHC stain should be able to assess IHC images according to markers’ expression levels and align with H&E models in the semantic space in the aspect of the tissue structure. With the limited publicly available IHC data, we exploit the trained H&E foundation model and introduce the “cross-stain transferring” pretext task, to deepen comprehension of pathological images stained in a different way. Specifically, we align the IHC representation with the IHC-style transferred H&E representation via a transferer drawing on the adaptive instance normalization, AdaIN [19]. This alignment injects structural information that is readily accessible in H&E images, as well as preserving diagnostic information rooted in IHC images.
@@
Fig. 1: An overview of PathoDuet. Left: two pretext tasks, cross-scale positioning and cross-stain transferring, are designed to develop H&E and IHC models. Right: a series of downstream tasks, covering both H&E and IHC ones, are used to evaluate models’ performance in application.
Fig. 2: An overall performance visualization. Each task is named as training dataset, (special settings,) evaluating metric. H&E tasks are colored purple, and IHC ones are yellow.
@@
A pretext token mechanism is introduced to unify the two proposed tasks. Both tasks require an auxiliary input in a different form, i.e., a much smaller patch or a staining hint. Contrary to designing a separate network to handle the additional input, an extra token with the auxiliary information is fed into the Vision Transformer (ViT) model [14] throughout the network training, and thereby combining the cross-scale or cross-stain information with original representation. The relation is subsequently exploited by a delicately designed module, termed a task raiser, to explicitly associate the two forms. This mechanism enriches the model’s capacity to discover and leverage intrinsic correlation between tasks and staining modalities in a lightweight way.
@@
The whole framework and ensuing models are collectively denoted as PathoDuet as shown in Fig. 1. This reflects the dual functionality of the framework, i.e., it offers two distinct strategies for developing foundation models: 1) by exploiting a shifted view of scales to discover broader semantic space, e.g., comprehending patches further from surrounding regions, 2) by progressing on the basis of other closely related and already exploited modalities, e.g., learning representation of IHC images from H&E model. The efficacy of our proposed methods is validated across a wide range of downstream tasks, covering both H&E and IHC images. We use a spider chart in Fig. 2 to visualize the overall performance across different tasks compared with powerful pathological models. These tasks range from classifying cancer tissue types at the patch level, categorizing WSIs, identifying cancer cells within IHC images, assessing the expression levels of IHC markers, to WSI-level IHC qualitative analysis. Our contributions are summarized as:
  • 1.
@@ 2.1 Self-supervised learning
Pretext tasks can be various. Some tasks aim at predicting properties of images, like rotation angles [15], flipping [47], etc. Some design small problems that help to learn features of images, like jigsaw puzzle solving [39] and relative position prediction of sub-regions [13]. Some exploit the process of image generation, including image colorization [58], super-resolution [29], inpainting [41], and exploiting generative adversarial networks (GANs) [62]. Besides, He et al. propose masked autoencoder (MAE) [17] that leads recent research into masked image modeling.
Among all these pretext tasks, similarity-based CL task has demonstrated their superiority, because they focus on invariant features instead of covariant ones. CL-based methods pull together the latent representations of similar images while keeping those of dissimilar images away from each other. These methods often regard different augmented views from the same image as a similar (positive) pair, and those from different images as dissimilar (negative) pairs. Therefore, CL-based methods typically contain two network branches. MoCo v1 [18] is a basic one with a symmetric structure, a momentum-update mechanism, a memory bank, and an InfoNCE loss that considers positive samples and negative samples together. SimCLR [7] adds a simple multi-layer perceptron (MLP) called projector after each backbone encoder and updates two branches concurrently without a memory bank. BYOL [16] further adds another MLP, predictor, after one of the branches (the online branch), while updating the other target branch with momentum, and changing the loss to cosine similarity loss which does not involve negative samples. SimSiam [8] proposes a similar structure of BYOL but in a symmetric way with its gradient stop technique to avoid collapse. In MoCo v3 [9], it adopts the BYOL’s structure but continues using the InfoNCE loss. DINO [5] and its later version DINO v2 [40] integrate previous works and achieve another state-of-the-art performance, thereby attracting applications to other fields.
@@ 2.2 Self-supervised learning in digital pathology
With the development of SSL frameworks in computer vision, the concept and some existing methods have been migrated to histopathology. Besides traditional pretext tasks, some works design specific pretext tasks in digital pathology. Magnification prediction [43, 26] and stain prediction [26] are two simple tasks that exploit the characteristics of pathological images. Resolution sequence prediction [47] is a dedicated task with easy supervision and remarkable performance. Others leverage the stain. RestainNet [60], as its name suggests, simulates the process of de-staining and restaining via separated Hematoxylin and Eosin channels. Ling et al. [33] use self-supervised methods to realize arbitrary stain transfer of pathological images from different domains. These methods though take into consideration the features of histopathological images, their performance depends on the association between the pretext tasks and the downstream tasks.
Some works are based on CL frameworks. Huang et al. [21], Li et al. [30], Ciga et al. [10], Kawai et al. [25] directly migrate traditional CL methods to medical data, while some studies take into account the uniqueness of histopathological images and propose a modified SSL framework. CTransPath [53] keeps the body of MoCo v3 but adds a pseudo positive selection mechanism to avoid false penalty of semantic similar patches and uses a hybrid CNN-Transformer as its backbone encoder. Meanwhile, Wang et al. also propose a clustering-guided contrastive learning method and the produced model RetCCL [52]. CS-CO [57] separates histopathological images into H-channel and E-channel and uses a cross-stain prediction task in the first phase and a CL-based method in the second phase. Abbet et al. [1] take advantage of domain information. Meanwhile, a branch of works has focused on the gigapixel WSIs [50, 51, 28, 2, 44, 55, 31]. Recently, pathological models pretrained with ultra-large amounts of data have shown their superiority [6, 12, 49]. Chen et al. [6] collect 100,000 slides to pretrain UNI with DINO v2 framework, and Virchow [49] with 1.5 million slides. Besides, some works look into multi-modal methods like visual-language learning and pretrain models on large-scale image-text datasets [20, 42, 36, 56], and the focus is more on leveraging strong language models.
@@ 3.1 Pretext token empowered SSL framework
The basic framework is based on a typical contrastive learning method, i.e., MoCo v3 [9]. The difference starts from the input. To mimic pathologists, we need an extra input to bring the information from a different scale or stain. A typical practice to handle two input types is employing two networks respectively with a following interaction module. However, we aim at learning a single powerful encoder for the downstream tasks. Therefore, we propose the pretext token mechanism to uncover the relation within the same encoder, thus deepening understanding of pathological images like pathologists.
The mechanism starts from adding a pretext token, which is concatenated with the partitioned mini-patches’ embeddings before being fed into the ViT. The token interacts with these embeddings and becomes an abstract association between the input pair in the attention blocks. After that, some simple and lightweight modules, termed as task raisers, utilize the association to perform pretext tasks. As illustrated in Fig. 3, the task raiser is instantiated as a positioner to position a local patch from its surrounding region in the cross-scale positioning task, or a transferer to generate IHC-style features from H&E images in the cross-stain transferring task.
In the following two subsections, we introduce in detail two pretext tasks we propose within this framework, and how pretext token and instantiated task raisers take effect in extracting information from two inputs. If not specified, we use the patch for the cropped image from a WSI and mini-patch for the output of patch embedding in ViT. Meanwhile, if the network takes only one kind of input, e.g., patches in patch network and IHC images in IHC network, the final input is formulated as $[\mathbf{x};\epsilon]$, where $\mathbf{x}=[x_{1};\dots;x_{L}]\in\mathbb{R}^{L\times C}$ represents the original input after patch embedding module in ViT. $\epsilon\in\mathbb{R}^{C}$ is a set of learnable parameters to hold the place for the pretext token so that the architecture of networks can be consistent and the usage of our models can be similar to that of a normal ViT. $L$ is the embedding length, and $C$ is the number of channels. When the network takes two inputs, the input changes to $[\mathbf{x}^{o};x^{p}]$, where $\mathbf{x}^{o}=[x^{o}_{1};\dots;x^{o}_{L}]\in\mathbb{R}^{L\times C}$ stands for the embeddings of the original input, e.g., a region or an H&E image, and $x^{p}\in\mathbb{R}^{C}$ for those of the pretext token input, e.g., a patch or an IHC crop.